PROPOSED CHANGES TO BY-LAWS OF INTERNATIONAL CONFERENCES ON AUTOMATED DEDUCTION

David A. Plaisted July 14, 1996

These proposed bylaws have been developed by David Plaisted in consultation with many people, including the Trustees. They specify a fully democratic system for the administration of CADE. Every detail of these bylaws has been designed to make CADE as democratic as possible and to give the membership a maximum voice in the running of CADE. The Trustees have also drawn up an alternative, partially democratic set of bylaws. A vote to decide which set of bylaws to adopt and whether to change CADE to a fully democratic system is scheduled for CADE 13. About 40 people have expressed their desire to see a fully democratic system adopted for CADE, and nine individuals including myself have specifically endorsed this final version. We encourage your support also.

This document lists places where the proposed bylaws differ from the current ones, accompanied by commentary about each change.

Changes have been indicated in italics. Italics may also indicate that nearby text has been deleted. Changes to the most recent version are indicated in small capitals.

ARTICLE III. MEETING OF MEMBERS MODIFIED TEXT (italicized text added):

Section 1. Meeting. There will be a membership meeting at each conference sponsored by the corporation. This will also be referred to as a Business Meeting in the sequel. At the membership meeting, business matters may be discussed, elections may be held, and motions may be made by the constituency and voted on by majority vote (a majority of the votes cast) of the constituency. The Trustees shall choose one of themselves, or the Program Chair, to have charge of the meeting. Motions passed at the Business Meeting or in a Special Election shall be binding on the Trustees. However, motions passed at the Business Meeting can be reversed by a vote in a Special Election.

Comment: This specifies more clearly what happens at the CADE business meeting. Motions passed are binding on the Trustees, but can be reversed in a special election.

ARTICLE IV. BOARD OF TRUSTEES OLD TEXT

Section 2. Number, Qualification and Tenure. The Board of Trustees shall consist of the Program Chairs of the last 4 CADE conferences, the Program Chairs of any CADE forthcoming conferences, the Secretary, the Treasurer, and up to three co-opted members. The co-opted members will be selected by the Board of Trustees and will normally serve for four years. Nothing in these By-Laws shall be construed to preclude the election of any current or former Trustee for any new term consistent with this By-Law.

NEW TEXT

Section 2. Number, Qualification, Election, and Tenure.

Subsection 2.1 Terms of Trustees; Three-Term Rule There will be nine Trustees. The terms of office of the Trustees begin at the conclusion of the CADE conference at which they are elected. The Trustees shall have terms of office that last for three CADE conferences, that is, beginning at the conclusion of the n-th CADE conference at which they are elected and ending at the conclusion of the (n+3)-th CADE conference. Three Trustees will be elected at each CADE conference. No Trustee may serve three successive terms of office. After serving two consecutive terms office, ending in calendar year n, a Trustee may not run for re-election again until calendar year n+2. However, a Trustee may serve three or more terms of office as long as no three terms of office are consecutive.

Subsection 2.2 Nominations of Trustees The Trustees serving at the time of the Business Meeting will select at least five individuals who, if they agree, will become nominees for the new Trustee positions; these may include the current Program Chair and incumbent Trustees who are eligible to serve again. These nominees will (if they consent) all become Candidates. Nominations for the Trustee positions may also be made by members from the floor at the Business Meeting, or by mail, electronic mail or fax to be received by the Trustees at least two weeks in advance of the Business Meeting. Two members are required to make a nomination before the Business Meeting. However, only one member suffices to make a nomination during the Business Meeting. Each member may only make one nomination (other than the nominations made jointly by the Trustees) before the Business Meeting. Individuals nominated in accordance with these conditions, who give their permission, also become Candidates.

Subsection 2.3 Election of Trustees New Trustees shall be elected at the CADE Business Meeting by a secret ballot. Each voter will be able to make a number of votes equal to the number of vacant positions (ordinarily, three). Suppose that there are n vacant Trustee positions to fill. If there are more than 2n+1 (twice n plus one) Candidates, the election will be in two rounds. In the first round, the 2n (twice n) Candidates receiving the largest number of votes will be selected and from these, the n Candidates receiving the largest number of votes in the second round will be chosen to be Trustees. The second round may be omitted if a motion to omit the second round is made and approved by the members after the first round. This motion need not be voted on by secret ballot. If there are 2n+1 or less Candidates, then the n Candidates receiving the largest number of votes in the first round will become Trustees. In the case of ties, tie-breaking decisions shall be made by vote of the Trustees.

Subsection 2.4 Vacancies Trustee positions which become unexpectedly vacant due to resignation, death, or for other reasons, may be filled at the election taking place at a following CADE conference, or in a Special Election.

Comment: This new text specifies the procedures for electing Trustees. There will be three Trustees elected each year for three year terms. The election will be by a two round system, where people cast a number of votes equal to the number of positions open. Limitations on successive terms of office are specified. Incumbent Trustees and current program chairs may be nominated by the Trustees, but do not become Trustees unless they are elected. The Trustees will make additional nominations for these positions.

MODIFIED TEXT

Section 9. Powers and Obligations of the Board of Trustees.

b) Select Program Chairs and Local Arrangement Chairs for future conferences.

Comment: Here we omitted the following text: “The Program Chairs will join the Board of Trustees.” In the proposed bylaws, the program chairs will not join the Board of Trustees unless they are elected.

OMITTED TEXT

The old items e) and f) under powers and obligations of the trustees are eliminated. These old items read as follows:

e) Co-opt up to three people as Trustees to serve on the Board of Trustees.

f) If there are additional vacancies on the Board of Trustees, elect by a majority affirmative vote person or persons to fill such unexpired terms.

Comment: In the proposed bylaws, all Trustees must be elected.

ADDED TEXT

Section 10. Special Elections. The Trustees may at times choose to hold Special Elections. A Special Election is defined as an election held between CADE conferences called to fill unexpected Trustee vacancies or decide issues relevant to CADE. It is not the same as a meeting of members. For Special Elections, the constituency shall be the members of the AAR with electronic mail access. The Trustees may also include AAR members without electronic mail access in the constituency, if they so choose. A Special Election may be conducted by electronic mail and/or postal mail. The election shall be by secret ballot, and decisions shall be made based on the relative numbers of votes cast for or against a proposal, or for various Candidates. Ties may be broken by a vote of the Trustees. If a Special Election is called to fill one or more unexpected Trustee vacancies, then a mechanism shall be provided for members to make nominations. It is permissible for such a Special Election to have two rounds.

Section 11. The Transitional Period. These bylaws shall go into effect 90 days after the close of the 1996 CADE conference. The Trustees (and ex officio Trustees) in office at that time shall be divided into a group of three whose terms of office expire at CADE 14, a group of three whose terms of office expire at CADE 15, and a group of three whose terms of office expire at CADE 16; any remaining Trustees’ terms of office will expire when these bylaws go into effect. The intention is to keep the number of Trustees at nine. The Trustees shall decide which of themselves belong to which of these groups. When these bylaws go into effect, three Trustees shall be elected at each CADE conference. All Trustees in office when these bylaws go into effect, and whose terms of office expire at CADE 14, will serve as non-voting Advisors to the Trustees for one year after CADE 14. By “non-voting” is meant that the Advisors do not vote at Trustee meetings (unless they are serving a term of office as a Trustee). Trustees whose terms of office expire at CADE 14, CADE 15, or CADE 16 may run for election when these bylaws go into effect. If their election occurs during or at the expiration of one of their terms of office as a Trustee, then they shall be considered as entering their second term of office, for purposes of the three-term rule. Otherwise, they shall be considered as entering their first term of office. In general, Trustees in office when these bylaws go into effect are subject to the three-term rule.

Comment: These sections specify what the headings imply, procedures for special elections and for the transitional period. A special election has the advantage of reaching a wider constituency than the attendees at a single CADE conference.

ARTICLE IX. AMENDMENTS OLD TEXT

Section 1. Power of Members to Amend By-Laws. The By-Laws of this corporation may be amended, repealed, or added to, or new By-Laws may be adopted by the vote or written assent of a majority of a quorum at a meeting duly called for the purpose according to the Articles or By-Laws.

Section 2. Power of Trustees to Amend By-Laws. Subject to the limitations of the Articles of Incorporation, these By-Laws, the General Non-Profit Corporations Law of Illinois, concerning corporate action that must be authorized or approved by the members of the corporation, the By-Laws of this corporation may be amended, repealed, or added to, or new By-Laws may be adopted by a resolution of the Board of Trustees.

NEW TEXT

Section 1. Power of Members to Amend By-Laws. The By-Laws of this corporation may be amended, repealed, or added to, or new By-Laws may be adopted by the vote or written assent of a majority of a quorum at a meeting duly called for the purpose according to the Articles or By-Laws.

Section 2. Amending By-Laws at a Special Election. The By-Laws of this corporation may also be amended, repealed, or added to, or new By-Laws may be adopted by secret ballot by majority vote (majority of the votes cast) at a Special Election as described in Article IV, Section 10 of these By-Laws.

Section 3. Restrictions on Amendment of By-Laws The By-Laws of this corporation may only be amended, repealed, or added to, or new By-Laws adopted, as specified in the preceding two Sections, except as required by applicable local, county, state, national, or international laws. However, the Trustees may modify the address of this corporation as given in Article I, Section 1 at any time, without a vote of the members.

Comment: In the proposed system, the Trustees may not amend the bylaws. This may only be done by a vote of the members. A simple majority suffices for this.

*********************************************************************

MINOR CHANGES

ARTICLE III. MEETING OF MEMBERS MODIFIED TEXT (italicized text added):

Section 4. Proxies. At any meeting of members, a member entitled to vote may vote by proxy executed in writing by the member or by his or her duly authorized attorney in fact. No proxy shall be valid after sixty (60) days from the date of its execution, unless otherwise provided in the proxy. A given member may serve as the attorney in fact, and cast a vote, for at most one member.

Comment: The restriction on proxies is intended to prevent one person from having so many proxies as to be able to control an election.

ARTICLE IV. BOARD OF TRUSTEES MODIFIED TEXT (italicized text added):

Section 4. Special Meetings. A special meeting of the Board of Trustees may be called by any three Trustees. The Trustees calling a special meeting of the Board may fix any place, either within or without the State of Illinois as the place for holding any special meeting of the Board called by them. The Trustees may hold meetings via electronic mail or in person, as they choose.

Comment: This clarifies that Trustee meetings may be held by electronic mail.

ARTICLE V. OFFICERS OLD TEXT

Section 2. Election and Term of Office. The Secretary and the Treasurer of the corporation shall be selected by the Board of Trustees. The President and Vice President shall be elected by and from the Board of Trustees. New offices may be created and filled at any meeting of the Board of Trustees. Each officer shall hold office until his or her successor shall have been duly selected and shall have been qualified unless specifically removed for cause. A person can hold more than one office, but will only be entitled to one vote.

NEW TEXT

Section 2. Election and Term of Office. The Secretary and the Treasurer of the corporation shall be selected by the Board of Trustees. The President and Vice President shall be elected by and from the Board of Trustees. New offices may be created and filled at any meeting of the Board of Trustees. The terms of the officers shall be for three CADE conferences, from the end of the n-th CADE conference to the end of the (n+3)-th CADE conference. Each officer shall hold office until his or her term of office has expired unless specifically removed for cause. Officers may serve for any number of consecutive terms of office, except that the President and Vice President may only serve during their elected terms of office. Only Trustees shall be entitled to vote at Trustee meetings. Officers are not considered as Trustees by virtue of their office. A member may hold more than one office. Trustees may also hold offices.

Comment: The terms of officers are specified to be three years, and officers who are not Trustees may not vote at Trustee meetings. The Trustees may still appoint a secretary but he/she may not vote unless he/she is elected as a Trustee.

OLD TEXT

Section 3. Removal. Any officer selected by the Board of Trustees may be removed by the Board of Trustees whenever in its judgment the best interests of the corporation would be served thereby, but such removal shall be without prejudice to the contract rights, if any, of the officer so removed.

NEW TEXT Section 3. Removal. The Trustees may vote at Trustee meetings to remove the President, Vice Presi-

dent, Secretary, or Treasurer from office at any time if, in their judgment, the best interests of the corporation shall be served by such removal. For this, five votes (more than half of the nine Trustees) in favor of removal are needed. Such removal shall be without prejudice to the contract rights, if any, of the officer so removed.

Comment: Procedures for removal of officers are further specified and the President and Vice President are explicitly included.

ARTICLE VI. COMMITTEES MODIFIED TEXT (italicized text added):

Section 2. Conference Committee. The Conference Committee for each conference shall consist of its Program Chair, Local Arrangements Chair and the President, Secretary and Treasurer of the Corporation. The Program Chair shall choose a Program Committee for the conference. The Conference Committee shall advise the Program Chair in the choice of the Program Committee. The Program Committee, under the supervision of the Program Chair, shall decide which submitted papers are accepted and rejected for the conference. Each Conference Committee shall be responsible for the planning and execution of its conference. It is empowered to establish subcommittees and enter into contractual relations to satisfy this purpose. In particular, both the Program Chair and the Local Arrangements Chair will be responsible for managing the finances associated with their part of the conference organization. They must keep accounts of these finances and these accounts must be approved by the Treasurer. The Treasurer will advance appropriate amounts of CADE Inc monies to the conference accounts and any profits made by the conference must be returned to CADE Inc at the end of the conference.

Comment: The duties of the Program Committee are further specified. This is unspecified in the existing bylaws, leaving theoretically open the possibility that there would be no Program Committee at all.

MODIFIED TEXT

Some text in ARTICLE IV. BOARD OF TRUSTEES, Section 5 was also modified to change the term “computer network” to “electronic mail.”